A uniform 3.0 kg disk of radius 30 cm is rotating around its central axis at an angular speed of 72 rad/s. At time t = 0, a man begins to slow it at a uniform rate until it stops at t = 20 s. (a) Find its angular acceleration. (b) By time t- 10 s, how much work had the man done? (c) For full 20 s, at what average rate did the man do work?

1
Expert's answer
2022-04-07T18:25:41-0400

"I=\frac 12 mr^2=0.135~kg\cdot m^2,"

"E=\frac{I\omega^2}2=350~J,"

"\varepsilon=\frac{\omega}{t_{20}}=3.6~\frac{rad}{s^2},"

"\omega_{10}=\omega-\varepsilon t_{10}=36~\frac{rad}s,"

"E_{10}=\frac{I\omega_{10}^2}2=87.5~J,"

"W=E-E_{10}=263~J,"

"P=\frac E{t_{20}}=17.5~W."